Left Direct Gregorian Technical Observing with IRTC: UT 2009 October 27

Observer: JHill (LBTO), DMiller (Tucson), JEsguerra (LBTO)
Telescope Operator: ACeranski
Instrument Support: TShih
Telescope Support: JUrban
SW support: (Tucson)

Summary

Cirrus clouds at the beginning, then clearing later.

Seeing between 0.5 and 1 arcsec during the first half, so transformation data is vastly improved from last night.

Startup problems with OSS and LSS. Lost 50 minutes at beginning due to an overwritten configuration file for AGW in today's software changes for GCS/oacontrol.

Sky was photometric, but seeing got very poor (2 arcsec) after midnight.

The new pointing model from last night works quite well.

Preparations

Software guys have renamed all the mountain workstations. lbtmu42 >>> ops2

LBT STC is here on the mountain for a meeting this afternoon.

00:32 Test preset failed correctly as MCSPU was still powered off.

00:40 Open Vent Doors

Details

All Times in UT

Open

00:45 Open shutter doors - T=+3.2 degC D=-28 degC, Wind light, western half of sky covered with thin-medium cirrus.

00:50 resetting direct rotator with Aaron holding bypass button

Sky flats for the guider

/OldRepository/AGW_Data/091027/

Taking B (0) guider flats 003 - 053 500 ms

Taking V (1) guider flats 054 - 104 1000 ms

Taking r' (2) guider flats 105 - 155 3000 ms

Pointing Check

01:02 pointto M5_0746 gs=0 ACQUIRE PAM- - Adjust IE=13 CA=29

01:06 pointto M5_0746 gs=2 ACTIVE PA=INDEF=187.48 - star way too bright ??

01:08 pointto M5_0746 gs=4 ACTIVE PA=INDEF=187.48 - star way too bright ??

01:10 pointto M5_1526 gs=4 ACTIVE PA=INDEF=94.36 - still crazy bright, but we are able to guide on it. The problem is that the guide probe did not move off axis. The stage is apparently not getting commands from GCS.

01:18 stop and restart GCS

01:19 pointto M5_1526 gs=4 ACTIVE PA=INDEF=94.36 - GCS crashed. Invalid buffer: RPCERROR

01:20 pointto M5_1526 gs=4 ACTIVE PA=INDEF=94.36 - still didn't move AGW probe.

01:23 stop GCSL, stopAGW, startAGW, home, restart GCS.

01:26 pointto M5_1526 gs=4 ACTIVE PA=INDEF=94.36 - probe moved (no, maybe only at park position), but acquisition grabbed the bright lines at bottom of image. Star may be too faint through the cirrus clouds.

Aaron applies global offsets in X and Y to M1/M2 to get M1 off the limit.

01:29 pointto StoneOagw gs=0 TelM="ACQUIRE" PAM-

OSS is hung up as last night. OSS ignores PSF requests for collimation. This time we have Paul nearby. Paul provides his professional opinion that OSS is hung completely. Then he kills and restarts it. Lost 20 minutes.

01:41 pointto M5_1526 ACQUIRE gs=0 PAM- Probe didn't move again. Events not displayed by IIF GUI ??

Clear Secondary Offload Corrections.

01:48 pointto M5_1526 ACQUIRE gs=0 PAM-

01:50 Events on IIF GUI appear just now from 10 minutes ago. Paul stops and restarts LSS.

01:52 pointto M5_1526 ACQUIRE gs=0 PAM- - Now events display promptly on IIFGUI.

Torsten discovers that configuration file was accidentally overwritten in the new install this morning. So the presets were moving AGW3 probe rather than AGW4. Total time lost to this problem was 50 minutes, but it was overlapping with the OSS problem and the LSS problem.

01:58 pointto M5_1526 ACQUIRE gs=0 PAM- Adjust IE=3 CA=17.

02:02 pointto M5_1526 gs=4 ACTIVE PA=INDEF=94.36

Clear C00 and collimation converges to 0.7 arcsec.

StoneO Transformation Data

02:08 pointto StoneOagw gs=0 TelM="ACQUIRE" PAM- IE=12 CA=17

Remove the Xglobal offset, so now we are at net +1 mm Y for M1 and M2.

02:13 pointto StoneOagw gs=1 TelM="ACTIVE" PA=INDEF=299.36 - seeing is 0.7 arcsec.

Doug puts the GCS guide scale factor back down to 0.75.

02:16 cl < home$tran_StoneO_00.cl
transtar (21) irtc.20091027.021820.fits cube 004
transtar (17)
transtar (11)
transtar (32)
transtar (9)
transtar (19)
transtar (4) guidestar is elongated

Up until here the IRTC was in the narrow field camera! Aaargh.

01:36 Change back to the IRTC wide field and J filter.

This data is taken through fairly thick cirrus clouds, but the seeing is in the 0.7-1 arcsec range on the guider.

This is the transformation we are using from oacontrol.conf.default dated 26-Oct 18:45 UT. (This is not the updated version that Jesper calculated from StoneA.)
        # sfp to agw transformation matrix
        sfpagw11=-1.001857726
        sfpagw12=1.6785432489e-2
        sfpagw21=-1.6785432489e-2
        sfpagw22=-1.001857726

The RA proper motion on the central star of StoneO has been corrected compared to last night.

02:39 cl < home$tran_StoneO_00.cl
transtar (21) irtc.20091027.024053.fits cube 011
transtar (17)
transtar (11)
transtar (32)
transtar (9)
transtar (19)
transtar (4)
transtar (8)
transtar (1)
transtar (37) no star found in acq, GCS grabbed noise in corner. Star out of position in IRTC.
transtar (21)
transtar (17)
transtar (11) irtc.20091027.031047.fits cube 023

Dewpoint is -34.6 degC - no wonder my lips are dry.

Seeing on the guider is 0.5 to 0.7 arcsec for this set. Still working through broken cirrus.

03:12 cl < home$tran_StoneO_90.cl
transtar (6) irtc.20091027.031342.fits cube 24
transtar (31)
transtar (20)
transtar (30) GCS grabbed noise instead of the faint star. Star is out of position on IRTC.
transtar (36)
transtar (16)
transtar (34)
transtar (9)
transtar (32)
transtar (19)
transtar (24)
transtar (10) AGW says probe is out of range. Star is out of position on IRTC.
transtar (11) irtc.20091027.034257.fits cube 036

03:44 cl < home$tran_StoneO_180.cl
transtar (24) irtc.20091027.034536.fits cube 037
transtar (18)
transtar (12)
03:51 PSF is complaining about bad new influence forces, even though the total correction forces appear to be small.
transtar (28)
transtar (6)
transtar (31)
transtar (20)
transtar (23) Couldn't move probe error. Star is out of position.
transtar (3) Couldn't move probe error. Star is out of position.
transtar (15) Couldn't move probe error. Star is out of position.
transtar (27) Couldn't move probe error. Star is out of position.
transtar (33) Couldn't move probe error. Star is out of position.

Tue Oct 27 04:02:28.826 2009  Preset alert for left IRTC mode ACTIVE: couldn''t move probe to position (SFP)
 179.117, 85.7183, focus: 14.152 :A

04:10 stop GCSL, stopAGW, startAGW, home, restart GCS.

04:11 change IRTC file prefix to "trans" instead of "irtc" to avoid an overwriting problem in new /Repository/. See Issue #2413.

04:12 Clear C00 on SX primary to get rid of bad influence forces.

04:15 pointto StoneOagw gs=1 TelM="ACTIVE" PA=INDEF=299.36 - Seeing was 0.55 to 0.65. Doug's GCSGUI crashed when we sent this preset, even though it was a fresh GCS started a few minutes.

04:18 cl < home$tran_StoneO_270.cl
transtar (1) irtc.20091027.041924.fits
transtar (3)
transtar (4)
transtar (5) (binary - don't use)
transtar (7)
transtar (8)
transtar (14)
transtar (21)
transtar (25)
transtar (27)
transtar (33) irtc.20091027.044332.fits cube 059
transtar (15)
transtar (29)
transtar (38) (USNO w/o pm)
transtar (39) (USNO w/o pm)

StoneA - Dice5

04:55 pointto StoneAagw gs=12 ACTIVE - Seeing is 0.6 arcsec

05:00 pointto StoneAagw gs=12 TelM="ACTIVE" PA=0

Change file prefix to dice.

05:01 dice5 {7 7 DETXY 4 loops } irtc.20091027.050147.fits

05:08 IRTC freezes up Doug's computer when he transfers IRTC downtown. While this was stuck, we missed several of the dice5 images. Then things started working again when I brought the IRTC window back to the mountain.

ls: No match.
   Cube sequence obtained: Exit
...Done! (success)
   ### Loop 1 position 4
   irc OffsetPointing 0. -7. -7. DETXY MOUNT true ABS left
...Done! (success)
   Image obtained: irtc.20091027.050635.fits
...Done! (failed: see LOGs/takeseq_050630.log)
   ### Loop 1 position 5
   irc OffsetPointing 0. 7. 7. DETXY MOUNT true ABS left
...Done! (success)
...Done! (failed: see LOGs/takeseq_050807.log)

   ### Loop 2 position 1 ##########
   irc OffsetPointing 0. 0. 0. DETXY MOUNT true ABS left
...Done! (success)
...Done! (failed: see LOGs/takeseq_050835.log)
   ### Loop 2 position 2
   irc OffsetPointing 0. -7. 7. DETXY MOUNT true ABS left
...Done! (success)
...Done! (failed: see LOGs/takeseq_050852.log)
   ### Loop 2 position 3
   irc OffsetPointing 0. 7. -7. DETXY MOUNT true ABS left
...Done! (success)
   Image obtained: irtc.20091027.050920.fits
ls: No match.
   Cube sequence obtained: Exit
...Done! (success)
   ### Loop 2 position 4
   irc OffsetPointing 0. -7. -7. DETXY MOUNT true ABS left
...Done! (success)

file prefix back to trans. Seeing is 0.5 to 0.8 arcsec in this sequence. Sky is clear.

05:30 cl < home$tran_StoneA_00.cl
transtar (12) irtc.20091027.053135.fits cube 081
transtar (16)
transtar (17)
transtar (18)
transtar (14)
transtar (12)
transtar (6)
transtar (5)
transtar (3)
transtar (12)
transtar (11) irtc.20091027.055527.fits cube 91

Accumulated Secondary Offload is X=-0.583 Y=-1.703 Z=+0.081 RX=-19.24 RY=+4.87
This should be added to the global offsets for M1.

05:56 cl < home$tran_StoneA_90.cl
transtar (12) irtc.20091027.055837.fits cube 92 (satellite went through acq image)
transtar (13)
transtar (15)
transtar (19)
transtar (20)
transtar (21)
transtar (16)
transtar (17)
transtar (18)
transtar (14)
transtar (12)

John's GCSGUI crashed just before this.

06:23 cl < home$tran_StoneA_180.cl
transtar (13) irtc.20091027.062452.fits cube 103
transtar (15)
transtar (19)
transtar (20)
transtar (21)
transtar (9)
transtar (2)
transtar (13) irtc.20091027.064134.fits cube 110

Seeing degraded to >1 arcsec during this sequence.

06:42 cl < home$tran_StoneA_270.cl
transtar (1) irtc.20091027.064434.fits cube 111
transtar (2)
transtar (8)
transtar (4)
transtar (6)
transtar (1)
transtar (2)

07:00 pointto StoneAagw gs=10 TelM="ACQUIRE" PAM- - hung waiting for rotator at -85 deg. Aaron had to unwrap it manually. Adjust IE=4 CA=20 (the first time in several hours) Lost 4 minutes.

07:09 pointto StoneAagw gs=12 TelM="ACTIVE" PA=90 Seeing was varying 0.6 to 1.4 arcsec.

Seeing was highly variable during the following sequence. It got up to 2 arcsec by Loop 5.

07:13 dice5
(off1   =                   7.) Offset1 in arcsec
(off2   =                   7.) Offset2 in arcsec
(coord  =                DETXY) COORD: RADEC, AZALT, or DETXY?
(exptime=                   1.) IRTC Exposure time in seconds
(Nexpo  =                   60) Number of sequence exposures
(loopN  =                    6) Number of dice5 iterations
At PA=90 Starting with irtc.20091027.071405.fits

   irc OffsetPointing 0. -7. -7. DETXY MOUNT true ABS left
...Done! (success)
   Image obtained: irtc.20091027.073512.fits
ls: No match.
   Cube sequence obtained: Exit
...Done! (success)
   ### Loop 3 position 5
   irc OffsetPointing 0. 7. 7. DETXY MOUNT true ABS left
...Done! (success)
Resmsg[0]  = GetImage command result: ERROR
Resmsg[1]  = [Error] IRSocket::GetImage::GetPar::res -> Could not read from socket.
   Image obtained: 1
ls: No match.
   Cube sequence obtained: Exit
...Done! (success)

08:04 added the global offsets reported above to the SXPMIRTCDCollimation.dat file.

08:07 Apply it and remove the previous secondary offload.

08:09 pointto StoneAagw gs=10 TelM="ACQUIRE" PAM-

08:11 pointto StoneAagw gs=12 TelM="ACTIVE" PA=0

Spherical Motion Test with OSS

08:13 Stop sending to PSF/PCS from GCS

08:15 Doug pushes move on sphere button and image dissappears. This is because it applied the two angle in an absolute way and wiped out the RX and RY values of our collimation.

08:18 Now Doug succeeds in RX/RY, but wipes out all the other X,Y,Z values.

The test fails, since the move on a sphere must operate in a relative fashion (adding on to the existing collimation), and not completely revise the hexapod position.

08:27 pointto StoneAagw gs=12 TelM="ACTIVE" PA=0 - Seeing is about 1 arcsec.

08:33 Doug appears to find a bug in move relative. The zero relative motion wiped out the existing RX / RY.

08:35 Repeat the previous failure.

08:36 Moving on a sphere zeros X only (if theta is non-zero), but Moving on a sphere zeros RX / RY (if theta is zero).

Elevation Discontinuities Test

08:50 pointto StoneAagw gs=10 TelM="ACQUIRE" PAM-

Initialize both M2 and M1 lookup tables.

Adjust *IE=6 CA=15 * at EL=36.2 deg.

08:55 pointto StoneAagw gs=10 TelM="ACQUIRE" PAM- Repeats within 1 sec.

08:56 elslewtohold 90 - mirror panic'd. Raise mirror. Clear C00.

09:06 Image comes back within 1 arcsec even after going to zenith and panicing mirror.

irtc.20091027.090828.fits

09:08 elslewtohold 90 and back to track. After collimating, star is within 2 arcsec of original spot.

09:13 elslewtohold 0 and back to track. Now the star has moved about 2 arcsec in Y, but is still within 3 arcsec of original spot.

09:17 pointto StoneAagw gs=10 TelM="ACQUIRE" PAM- We don't see any effects beyond the open-loop tracking drift.

09:18 elslewtohold 90. (stop drives at zenith for Jorge to make an accelerometer adjustment)

irtc.20091027.093746.fits Star is at same position.

Field Data on BS9122

09:39 pointto BS9122 gs=0 ACQUIRE PAM- Didn't find star in initial acquisition. IE=14 CA=56 at EL=61

09:47 pointto BS9122 gs=0 ACTIVE PA=0 - Seeing is 0.6 arcsec on Guider. Collimate to 150 RMS

9:50

10:00 pointto BS9122 gs=0 GUIDE

Collimated on-axis and then went to off-axis guide star and could not find it. wfsc 1-8

10:14 pointto M5_0314 TelMode=ACQUIRE gs=0 Star off to the side. Aaron set IE=2 CA=20

10:20 pointto BS9122 TelMode=GUIDE gs=0 PA=225 Central star near the top of the Acquisition Image.

10:24 pointto BS9122 TelMode=GUIDE gs=18 PA=221 Guide star near the top of the Acquisition Image, almost the same as the central star

10:24 pointto BS9122 TelMode=GUIDE gs=18 PA=221 IE=-17 CA=14

10:24 pointto BS9122 TelMode=GUIDE gs=0 PA=221 Collimate (see table below)

Persson's Star = BS9122
Start time = 10:30

Guide Star x y PA Doug's Angle wfscimage000*
0 -0.028 0.115 221 -347 9-12
18 -1.28 89.84 221 -350 13-15
0 0.0 0.0 0.0 -351 16
17 0.67 118.579 -317. -445 17-19
0 0.10 0.01 312 -446 20-21
39 0.04 110.45 42.09 -538 22-24
0 0.1 0.01 342.09 -481 25-26
0 0.0 0.0 42.09 -183 27
42 -0.03 146.478 147.49 -290 28-30
0 0.0 0.0 147.99 -290 31-32
14 -0.069 156.828 225.62 -372 33-35
0 0.0 0.0 225.62 -375 36-37
11 0.016 158.765 296.79 -448 38-40
0 0.0 0.0 296.79 -451 41-42
24 0.029 145.585 349.12 -505 43-45
0 0.0 0.0 349.12   46-47
46 0.079 187.408 93 -254 48-50
0 0.0 0.0 93 -262 52-53
43 0.041 180.081 146.78 -319 54-56
0 0.0 0.0 146.78 -323 57-58
12 0.0 138.568 265.01   59-61
0 0.0 0.0 265.01 -447 62-63
36 -0.0 153.422 16.8 -202 64-66
0 0.0 0.0 16.8 -205 67-68
39 -0.02 110.509 42.09 -233 69-71
0 0.0 0.0 42.09 -236 72-73

11:39 Rotator went into hold mode. PresetTelescope waited and finally timed out after 300 second. IIF failed preset message did not say

12:09 Doug's GCS Gui in Tucson died

Vibration Data

12:18 pointto M5_0314 TelMode=ACTIVE gs=0 Oops, M5 is too bright

12:18 pointto M5_0314 TelMode=ACTIVE gs=0 Collimate 780 nm, 340 nm,

12:25 Test with 100 frames vibr: 148,149, 150

12:20 /Repository/20091027/irtc.20091027.123059.fits single image with header information

12:33 /Repository/20091027/irtc.20091027.123323.fits single image with header information

12:32 /Repository/20091027/vibr_00151_cube.fit exp=0.00005 s 1056Hz 3000:32x32 frame

12:36 /Repository/20091027/vibr_00152_cube.fit exp=0.00005 s 1056Hz 30000:144,120:32x32 frame

12:38 /Repository/20091027/irtc.20091027.123819.fits single image with header information

13:08 /Repository/20091027/vibr_00153_cube.fit exp=0.0001 s 200Hz 30000:144,120:32x32 frame

13:38 Background /Repository/20091027/vibr_00154_cube.fit exp=0.0015 s 200Hz 100:144,120:32x32 frame

13:39 Background /Repository/20091027/vibr_00155_cube.fit exp=0.0015 s 200Hz 1000:144,120:32x32 frame

13:40 Background /Repository/20091027/vibr_00156_cube.fit exp=0.0005 s 200Hz 1000:144,120:32x32 frame

13:41 Background /Repository/20091027/vibr_00157_cube.fit exp=0.0005 s 200Hz 100:144,120:32x32 frame

Close

-- DougMiller - 21 Oct 2009
I Attachment Action Size Date Who Comment
daq.20091027.gifgif daq.20091027.gif manage 43 K 28 Oct 2009 - 05:35 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r7 - 12 Nov 2009, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback